OpenStudy (anonymous):

Please help D: |3x - 4| - 8 = 6

OpenStudy (anonymous):

i got x=6

OpenStudy (anonymous):

|3x-4|=6+8 |3x-4|=14 Now, |3x-4| can either mean +(3x-4) or -(3x-4) So we have two cases, Incase, it is +(3x-4) then, 3x-4=14 3x=14+4 3x=18 x=18/3=6 Incase it is -(3x-4) then, -3x+4=14 Or, 4-14=3x -10/3=x